A family has four children. If X is a random variable that pertains to the number of female children, what are the possible values of X?


1
Expert's answer
2022-02-14T10:29:23-0500

The possible values of "X" are: 0, 1, 2, 3, 4


"P(X=0)=P(BBBB)=(\dfrac{1}{2})^4=\dfrac{1}{16}"

"P(X=1)=P(GBBB)+P(BGBB)"

"+P(BBGB)+P(BBBG)=4(\dfrac{1}{2})^4=\dfrac{1}{4}"

"P(X=2)=P(GGBB)+P(GBBG)"

"+P(GBGB)+P(BGGB)+P(BGBG)"

"+P(BBGG)=6(\dfrac{1}{2})^4=\dfrac{3}{8}"

"P(X=3)=P(GGGB)+P(GGBG)"

"+P(GBGG)+P(BGGG)=4(\dfrac{1}{2})^4=\dfrac{1}{4}"



"P(X=4)=P(GGGG)=(\dfrac{1}{2})^4=\dfrac{1}{16}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 & 3 & 4 \\ \hline
 p(x) & 1/16 & 1/4 & 3/8 & 1/4 & 1/16 \\
 
\end{array}"